# Ranking Generated Answers

### v1

@@ 5 Conclusions
We developed a method to scalably and automatically evaluate generated answers to consumer health questions. Our NRP score evaluates model effectiveness by ranking answers in comparison with human-written documents. NRP does not rely on ‘gold standard’ answers, but uses existing texts. We investigated the factors that led to higher effectiveness when using the measure, and whether human evaluations of generated answers agree with the rankings from an effective ranking model. However, the results of our experiments may be limited in their ability to discriminate highly effective LLMs as the dataset we used was potentially too easy for these models. We leave this investigation for future work. We demonstrate that state-of-the-art ranking functions, once validated using a test collection with multi-dimensional relevance assessments, can be used to effectively discriminate high-quality from low-quality responses. The ranking functions can even be used to discriminate between responses from similar LLMs. Furthermore, once the test collection and ranking function are finalized, other LLMs can be evaluated using the setup in an entirely offline and scaleable manner. The experiments in this paper focus on evaluating generated answers for consumer health search questions. For future work, we will develop and adapt additional test collections for other domains. Another avenue for future research is the extension of NRP to assess retrieval-augmented generation (RAG) systems that ground their answers in documents by referring to them. This paper contributes a new way to automatically assess LLM capabilities by evaluating them in the context of human-annotated documents.

### v2

@@ 6 Conclusions
We developed a method to scalably and automatically evaluate generated answers to open-ended questions. Our NRP score evaluates model effectiveness by ranking answers in comparison with human-written documents. NRP does not rely on ‘gold standard’ answers, but uses existing texts. We investigated the factors that led to higher effectiveness when using the measure, and whether human evaluations of generated answers agree with the rankings from an effective ranking model.
We demonstrate that state-of-the-art ranking functions, once validated using a test collection with multi-dimensional relevance assessments, can be used to effectively discriminate high-quality from low-quality answers. The ranking functions can even be used to discriminate between answers from similar LLMs with different prompting strategies. Furthermore, once the test collection and ranking function are finalized, other LLMs can be evaluated using the setup in an entirely offline and scaleable manner.
The experiments in this paper focused on evaluating generated answers for consumer health search questions on the CLEF 2021 eHealth dataset. For future work, we will develop and adapt additional test collections for other domains. Another avenue for future research is the extension of NRP to assess retrieval-augmented generation (RAG) systems that ground their answers in documents by referring to them. This paper contributes a new way to automatically assess LLM capabilities by evaluating them in the context of human-annotated documents.
Acknowledgements
This publication has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070014 (OpenWebSearch.EU, https://doi.org/10.3030/101070014).